The motion of a car is described by x(t)=2t – 0.4ť. Calculate the instantaneous velocity at t = 4 sec.
Answer #1

x(7) = position da x(Z) = 3+ - 0.473 nowvelocity (0) dx (7) dt dt. d 2-3x0.47? 10 = (27-01475) dt at t= a see; no Z-45 - 2- 3